# Tidal deformability in neutron stars from a microscopic point of view

### v1

@@ I Introduction
A fully microscopic equation of state (EoS) up to central densities of the most massive stars – potentially involving non-nucleonic degrees of freedom and phase transitions – is not within reach. Nevertheless, neutron stars are powerful natural laboratories for constraining theories of the nuclear EoS Abb17a; Abb17b; Abb18; Abb19; Mil19; Mil21 and studying matter under extreme conditions.
@@
A mix of all fundamental forces existing in nature is present in a neutron star, which continues to motivate the large interdisciplinary interest in its properties and the intense effort to constrain those properties both observationally and experimentally, with various levels of statistical and systematic uncertainties, see e.g. Ref. Soum+18. When working with microscopic predictions (as opposed to phenomenological models or statistical analyses), one must be mindful of the theory’s limitations and identify the best ways to extract and interpret information from the observational or terrestrial data.
In this paper, we focus primarely on the tidal properties of neutron stars, which reflect directly on the emitted gravitational wave signal, offering the opportunity of detecting the latter to constrain the former FR2012; BR2017; Bai2019. This prospect was first realized in August of 2017, with the detection of gravitational waves generated by a coalescing neutron star binary. This groundbreaking event was made possible by the second-generation ground-based detectors LIGO AL and Virgo Acern+2015, an event known as GW170817 Abb17a. We can expect future observations to detect the inspiral and coalescence of more compact binary systems.
@@
During a neutron star merger event, the gravitational field of one neutron star on the other induces tidal deformations, which are encoded in the late-inspiral gravitational wave signal. More precisely, when a neutron star is placed in the perturbing tidal gravitational field generated by its companion compact star, its shape will be distorted, resulting in an induced quadrupole moment. The induced quadrupole moment of the neutron star will impact the energy of the system and increase the rate of emission of gravitational waves, with the result that the binary system emits more energy and evolves faster towards the collision and merger FH2008; Hind2008.
Compact binaries that are detectable through gravitational waves are natural systems to observe this type of interaction, which opens the way to finding additional constraints to the neutron star EoS. A stiffer EoS, where the pressure increases rapidly with density, yields stars with larger radii and larger tidal effects. Therefore, tidal distortion reflects the stiffness of the EoSs, with the inspiral proceeding more rapidly for stars with larger tidal deformability.
@@
As we review below, for a given neutron star mass, the tidal deformability is strongly correlated with the star radius and, thus, the star compactness, M/R. Keeping in mind that accurate and precise neutron star radius measurements are challenging Khad+2025, the importance of tight constraints on global neutron star properties that depend on M/R is self-evident. We also recall that the radius of the canonical neutron star correlates with the (model dependent) slope of the symmetry energy around saturation density.
@@
This paper is organized as follows. In Sec. II, we review briefly our theoretical ingredients, omitting details that have been published elsewhere. In Sec. III, we construct high-density continuations of the microscopic EoS, paying particular attention to the speed of sound in stellar matter. In Sec. IV.1, we outline the calculation of the tidal deformability, which can be incorporated in the solution scheme used for the well-known Tolman-Oppenheimer-Volkoff (TOV) differential equations for $\frac{dP}{dr}$ and $\frac{dm}{dr}$. Our results, along with other predictions and constraints, are discussed in Sec. IV.2. Our main takeaways are summarized in Sec. V.
II Basic concepts in chiral effective field theory
The microscopic part of our equation of state for cold $\beta$-equilibrated neutron matter is based upon high-precision two-neutron potentials EMN17 and includes the chiral three-neutron forces that appear at the specified order. For a comprehensive review of nuclear forces based on chiral effective field theory (EFT), see Ref. Love.
Given an energy scale and degrees of freedom appropriate at that scale, an EFT comprises all interactions consistent with the symmetries that govern those degrees of freedom. For the nuclear problem, relevant degrees of freedom are pions (Goldstone bosons), nucleons, and $\Delta(1232)$ isobars. We use the delta-less chiral EFT. To begin with, one writes the most general Lagrangians describing all interactions between pions, nucleons, and pions with nucleons. Because pion interactions must vanish at zero momentum transfer and in the chiral limit, where the pion mass, $m_{\pi}$, goes to zero, the corresponding Lagrangian is expanded in powers of spatial derivatives or pion masses. From these Lagrangians, an infinite number of Feynman diagrams can be generated, which seems to render the theory unmanageable. The strategy is then to design a scheme for ordering the diagrams according to their importance – the essence of Chiral Perturbation Theory (ChPT). Nuclear potentials are defined by the irreducible types among these graphs. These graphs are then analyzed in terms of powers of $Q$, with $Q=p/\Lambda_{b}$, where $p$ is generic for a momentum, (nucleon three-momentum or pion four-momentum), or the pion mass, and $\Lambda_{b}\sim m_{\rho}\sim$ 0.7 GeV (with $m_{\rho}$ the mass of the $\rho$ meson) is the breakdown scale Fur15. Determining the power ($\nu$) has become known as power counting.
@@ II Basic concepts in chiral effective field theory
While the predictions at N2LO are fully ab initio, a warning is in place for current N3LO calculations. As pointed out in Ref. EKR20, there is a problem with the regularized 3NF at N3LO (and higher orders) in all present nuclear structure calculations. The N3LO 3NFs currently in use are regularized by a multiplicative regulator applied to the 3NF expressions as derived from dimensional regularization. This approach leads to violation of chiral symmetry at N3LO and destroys the consistency between two- and three-nucleon forces EKR20; Ep+22. Consequently, none of the current calculations that include 3NFs at N3LO (and beyond) can be considered truly ab initio. An appropriate symmetry-preserving regulator EKR20 should be applied to the 3NF at N3LO from Refs. Ber08; Ber11. At the present time, reliable predictions exist only at N2LO, NLO, and LO. However, for the few fully ab initio calculations, the precision at N2LO is unsatisfactory. A first step towards deriving consistently regularized nuclear interactions in chiral EFT has been proposed in Refs. KE23a; KE23b. It requires the cutoff to be introduced already at the level of the effective Lagrangian. A path integral approach KE23a can then be applied to the regularized chiral Lagrangian to derive nuclear forces through the standard power counting of chiral EFT.
Throughout the paper, we will show results at the (fully consistent) third order (N2LO), and at the highest order which we have considered (fourth order, or N3LO). For details on how our EoS are built, the reader is referred to Refs. SM21a; SM22.
@@ III.1 Subluminar and non-conformal
The content of this section is based on Ref. S+A_2025, where we discussed how causality and maximum-mass constraints pose considerable restrictions on the general features of the high-density EoS, if such continuation follows an ab intio portion.
Parametrizations of the high-density EoS in terms of the speed of sound can be used as an alternative to, or in combination with the popular polytropic extensions. Defining $\rho_{0}$, $\epsilon_{0}$, and $P_{0}$ as the baryon density, the energy density, and the pressure at threshold (the density at which the EoS parametrization is attached to the previous piece), we write Kan+21; Tew18, for $i\geq 1$,
@@
where the constants $c_{1}$ and $c_{2}$ are determined from continuity of the speed of sound and its derivative at the threshold density.
@@
We found S+A_2025 that a better ansatz, for the purpose of achieving high maximum masses while respecting causality at any density, is to first continue the EoS with a relatively steep polytrope,
$P(\rho)=\alpha(\rho)^{\gamma}\;,$ (6)
and then follow with a parametrization as in Eqs. (45), which will remain causal by construction. The matching densities are 0.277 $fm^{-3}$ and 0.563 $fm^{-3}$ S+A_2025. The resulting M(R) curves are shown on the LHS of Fig. 1 for both N3LO (blue) and N2LO (red). For the dashed curves, the first extension is a polytrope with $\gamma$ = 3.3, followed by pressure values obtained from Eq. (5) with the speed of sound (SoS) given in Eq. (4). The solid curves (same color convention) have been obtained with $\gamma$ = 3.8, a value beyond which the EoS begins to violate causality. On the RHS, we show the dimensionless speed of sound squared corresponding to the curves on the left. Table 1 displays the maximum mass, its radius, the central density, and the radius of the canonical mass neutron star, for the curves in Fig. 1. It is important to recall that the radius of a 1.4 $M_{\odot}$ is sensitive to the pressure at normal nuclear densities and thus it can pose constraints on ab initio theories of the EoS at those densities where the theories can be applied reliably.
Figure 1: Left: M(R) curves at fourth order (N3LO, blue) and at third order (N2LO, red) of ChPT. Dashed curves: the first extension is done using a polytrope with $\gamma$ = 3.3, followed by pressure values given by Eq. (5) together with Eq. (4); Solid curves: obtained with $\gamma$ = 3.8, a value beyond which the EoS begins to violate causality. Right: Dimensionless speed of sound squared corresponding to the curves on the left. Same color and pattern conventions. The figure is taken from Ref. S+A_2025.
In Ref. S+A_2025, we concluded that a polytrope connecting the chiral EFT predictions with a causality-maintaining parametrization has a limited range of powers. We underlined that this scenario is fundamentally related to the softness of the chiral predictions. Therefore, the nature of the predictions at normal density have a far-reaching impact, which extends to densities up to a few times normal density.
Table 1: Some neutron star properties corresponding to the red and the blue M(R) relations shown in Fig. 1.The Table is taken from Ref. S+A_2025.
@@ III.2 Subluminar and conformal
In the QCD limit of deconfined quarks in the presence of asymptotic freedom, quarks should behave like free fermions. Some perturbative QCD calculations BWK23 support the conformal limit, $(v_{s}/c)^{2}=1/3$. We observed that an EoS that’s subconformal at all densities (that is, Eq. (4) with the asymptotic limit replaced by 1/3), cannot generate sufficiently large masses. On the other hand, the speed of sound can be asymptotically conformal and non-monotone. A scenario where the speed of sound peaks around few to several times nuclear density and then falls back to approach the QCD limit for deconfined quark matter would signify some sort of phase transition with the conformal limit reached well beyond central densities of the heaviest observed neutron stars. The superconformality condition, $\Big(\frac{v_{s}}{c}\Big)^{2}>1/3$, satisfied on the average in neutron stars may have fundamental implications on the trace of the energy-momentum tensor at the center of rotating neutron stars Raissa+24.
Here, we will explore parametrizations where the speed of sound is consistent with the conformal limit as the density approaches infinity. This can be simulated adding a second, skewed gaussian function to Eq. (4), and of course changing the asymptotic value to 1/3:
@@
We solve for $c_{1}$ and $c_{2}$ as done above, matching Eq. (7) and its derivative to the values at threshold, within a chosen parameter space for the remaining three coefficients, $c_{3},c_{4},c_{5}$. The widths $w$ and $w_{2}$ are obtained by normalizing both gaussians in the standard way. Although this procedure can provide a very large number of possibilities, we observed that the roots, $c_{1}$ and $c_{2}$, converge easily within limited ranges for the free parameters: $c_{3}\approx[0,1]$, $c_{4}\approx[1,4]fm^{-3}$, $c_{5}\approx[-50,50]$. In fact, the roots tend to cluster around a few pairs of $c_{1},c_{2}$ values, weakly sensitive to the large variations in $c_{5}$, and only moderately sensitive to variations of $c_{3}$ and $c_{4}$. As a consequence, those parametrizations produce very similar M(R) results. A representative case is shown by the green curves on the left-hand side of Fig. 2, where we have chosen a polytropic index of 3.5 to ensure that the new curves would stand out in between the other two sets from Fig. 1. The corresponding curves for the dimensionless speed of sound are seen to approach the conformal limit on the right-hand side of Fig. 2. We also note that implementing the conformal limit at all densities would not allow sufficiently large masses. Maximum masses as large as those recently observed require a rapid rise of the pressure soon after the first matching point.
It is interesting to observe that the left panels in Fig. 1 and Fig. 2 are consistent with each other, in the sense that the green curves are what one would expect from the non-conformal calculation and an adiabatic index that’s in between the other two. In contrast, the green predictions on the right are dramatically different from the red or blue predictions. This suggests that the maximum mass is only weakly sensitive to the super-high density behavior of the EoS.
Figure 2: Left: the red and blue M(R) curves are as described in Fig. 1. Green curves: M(R) relations at N3LO (dashed) and at N2LO (solid) where the polytropic part of the EoS has adiabatic index equal to 3.5, and the last continuation is based on the conformal speed of sound parametrization, Eq. (7). Right: Dimensionless speed of sound squared corresponding to each of the cases displayed on the left. See text for additional explanation.

### v2

@@ I Introduction
A fully microscopic equation of state (EoS) up to central densities of the most massive stars – potentially involving non-nucleonic degrees of freedom and phase transitions – is not within reach. Nevertheless, neutron stars are powerful natural laboratories for constraining theories of the nuclear EoS Abbott and others (2017); Abbott, B. P. and others (2017, 2018); Abbott and others (2019); Miller and others (2019, 2021) and studying matter under extreme conditions.
@@
A mix of all fundamental forces existing in nature is present in a neutron star, which continues to motivate the large interdisciplinary interest in its properties and the intense effort to constrain those properties both observationally and experimentally, with various levels of statistical and systematic uncertainties, see e.g. Ref. Soumi et al. (2018). When working with microscopic predictions (as opposed to phenomenological models or statistical analyses), one must be mindful of the theory’s limitations and identify the best ways to extract and interpret information from the observational or terrestrial data.
With the first detection of gravitational waves (GW) on September 14, 2015 from LIGO, major accomplishments have been made within the fields of physics and astronomy Miller and Yunes (2019). Gravitational waves tied up one of the last loose ends in the theory of general relativity with remarkably high precision Cervantes-Cota and others (2016). Our present understanding of black holes and neutron stars also rely heavily on data collected from GW.
Just as black holes release immense gravitational waves, neutron stars are remarkable sources of these waves. Exotic processes are triggered within a neutron star’s core due to the enormous pressure, which can forcefully eject mass at relativistic speeds. Consequently, gravitational waves become vital means of measuring these fascinating astrophysical objects and the dynamic processes within them Lasky (2015). For gravitational waves to produce discernible signals, a crucial requirement is the presence of asymmetrical deformations in the neutron star. Alternative mechanisms, such as pulsar “glitches,” events where neutron stars undergo a sudden “jump” in spin frequency, or magnetar flares, where a neutron star with an exceptionally high magnetic dipole experiences magnetic decay, can also induce substantial gravitational waves Lasky (2015).
@@
During a neutron star merger event, the gravitational field of one neutron star on the other induces tidal deformations, which are encoded in the late-inspiral gravitational wave signal. More precisely, when a neutron star is placed in the perturbing tidal gravitational field generated by its companion compact star, its shape becomes distorted, resulting in an induced quadrupole moment. The induced quadrupole moment of the neutron star will impact the energy of the system and increase the rate of emission of gravitational waves, with the result that the binary system emits more energy and evolves faster towards the collision and merger Flanagan and Hinderer (2008); Hinderer (2008). Clearly, compact binaries that are detectable through gravitational waves are natural systems to observe this type of interaction, which opens the way to finding additional constraints to the neutron star EoS. A stiffer EoS, where the pressure increases rapidly with density, yields stars with larger radii and larger tidal effects. Therefore, tidal distortion reflects the stiffness (softness) of the EoSs, with the inspiral proceeding more rapidly for stars with larger tidal deformability.
In this paper, we focus primarily on the tidal properties of neutron stars, which have a direct impact on the emitted gravitational wave signal, offering the opportunity of detecting the latter to constrain the former Faber and Rasio (2012); Baiotti and Rezzolla (2017); Baiotti (2019). This prospect was first realized in August of 2017, with the detection of gravitational waves generated by a coalescing neutron star binary. This groundbreaking event, known as GW170817 Abbott and others (2017), was made possible by the second-generation ground-based detectors LIGO Adams and others (2015) and Virgo Acernese and others (2015). We can expect future observations to detect the inspiral and coalescence of more compact binary systems.
@@
As we review below, for a given neutron star mass, the tidal deformability is strongly correlated with the star radius and, thus, the star compactness, C, defined as M/R. Keeping in mind that accurate and precise neutron star radius measurements are challenging Khadkikar and et al. (2025), the importance of tight constraints on global neutron star properties that depend on M/R is self-evident. We also recall that the radius of the canonical neutron star correlates with the (model dependent) slope of the symmetry energy around saturation density.
Although this paper could be seen as an extension of Ref. Sammarruca and Ajagbonna (2025), we note that it is an important extension. The tidal deformability encodes the impact of the EoS on the phase evolution of the waveform, which we have not previously confronted with our EoS. Tidal deformation is yet another astrophysically observable macroscopic property of neutron stars that can be used to study their interiors. Currently, Bayesian statistical analyses are very popular in the nuclear physics community, as they should be. Our approach is not a statistical analysis, in that we make predictions remaining as close as possible to the chiral theory of nuclear forces, and follow new constraints as they become available. We believe that our findings in Ref. Sammarruca and Ajagbonna (2025) are valuable and should be tested against these independent data. Furthermore, while this paper focuses on a particular, fundamentally important physical property of neutron stars, we wish to place it within our ongoing microscopic studies of the nuclear equation of state, where we keep a consistent path from free-space nucleon-nucleon ($NN$) data to neutron-rich nuclei and to neutron stars, as illustrated in Fig. 1. “MBT” stands for many-body theory, while SNM, NM, and NS signify symmetric nuclear matter, neutron matter, and neutron-star matter, respectively.
Figure 1: (Color online) Schematic representation of our ab initio pipeline.
@@
This paper is organized as follows. In Sec. II.1, we review briefly our theoretical ingredients, omitting details that have been published elsewhere. In Sec. III, we construct high-density continuations of the microscopic EoS, paying particular attention to the speed of sound in stellar matter. In Sec. IV.1, we outline the calculation of the tidal deformability, which can be easily incorporated in the solution scheme used for the well-known Tolman-Oppenheimer-Volkoff (TOV) differential equations for $\frac{dP}{dr}$ and $\frac{dm}{dr}$. Our results, along with other predictions and constraints, are discussed in Sec. IV.2. Our main takeaways are summarized in Sec. V.
II Theoretical tools
II.1 Basic concepts in chiral effective field theory
The first step in a development of an EFT is the identification of a “soft scale” and a “hard scale.” For this purpose, guidance can be found in the hadron spectrum, where a large separation exists between the mass of the pion and the mass of the vector meson $\rho$. It is therefore natural to identify the pion mass as the soft scale while the mass of the $\rho$ sets the hard scale, approximately 1 GeV. Moreover, since quarks and gluons are ineffective degrees of freedom in the low-energy regime, pions and nucleons are the appropriate degrees of freedom of the EFT. The connection between QCD and the EFT is established through the symmetries of low-energy QCD. At this point, we can write the most general Lagrangian consistent with those symmetries (and their breaking). Following the prescription of the theory as reviewed in Ref. Machleidt and Entem (2011), the QCD Lagrangian is given by:
$\mathcal{L}=\bar{q}(i\gamma^{\mu}\mathcal{D}_{\mu}-\mathcal{M})q-\frac{1}{4}\;\mathcal{G}_{\mu\nu,a}\mathcal{G}^{\mu\nu}_{a}\;,$ (1)
where ‘$q$’ is the quark field, $\mathcal{D}_{\mu}$ represents the gauge covariant derivative, $\mathcal{M}$ is the quark mass matrix, and $\mathcal{G}^{\mu\nu}_{a}$ is the gluon strength field tensor.
Chiral symmetry is conservation of “handedness,” and is an exact symmetry for massless particles. Chiral symmetry occurs in the limit of vanishing quark masses, which amounts to dropping the quark mass matrix term in the above Lagrangian. In fact, such term is responsible for explicit breaking of chiral symmetry, as can be seen from the following. The quark mass matrix,
$\mathcal{M}=\begin{pmatrix}m_{u}&0\\ 0&m_{d}\end{pmatrix}$ (2)
can be recast in terms of the identity matrix and the third Pauli spin matrix:
$\mathcal{M}=\frac{(m_{u}+m_{d})}{2}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}+\frac{m_{u}-m_{d}}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\;$. (3)
Clearly, the first term respects isospin symmetry while the second term vanishes if the masses of the “up” and “down” quarks are equal. Thus, the small difference in the quark masses breaks isospin symmetry. On the other hand, the expression above breaks chiral symmetry explicitly as a result of the non-zero quark masses. However, since the masses of the “uu” and “dd” quarks are very small compared to typical hadronic masses, explicit breaking of chiral symmetry is a small effect.
Next, we briefly address the spontaneous breaking of chiral symmetry, for which there is clear evidence in the hadron spectrum. The spontaneous breaking of a global (as opposed to local) symmetry is accompanied by the appearance of a so-called massless “Goldstone Boson.”The particle which fulfills these requirements is the pion, an isospin triplet pseudoscalar boson. The pion is light relatively to the other mesons in the hadron spectrum but not massless, which is due to the explicit chiral symmetry breaking from the non-vanishing quark masses.
Having identified pions and nucleons as the appropriate degrees of freedom, one can then proceed to construct the Lagrangian of the effective theory:
$\mathcal{L}_{eff}=\mathcal{L}_{\pi\pi}+\mathcal{L}_{\pi N}+\mathcal{L}_{NN}+$... (4)
This effective Lagrangian is expanded in the form of the “soft scale” over the “hard scale”, $\frac{Q}{\Lambda_{\chi}}$. $Q$ is of the order of the pion mass, whereas $\Lambda_{\chi}$ is the energy scale of chiral symmetry breaking, approximately 1 GeV.
Through a scheme known as power counting, the most important contributions to the effective Lagrangian are accounted for first, with increasing order resulting in consistently smaller terms. While the expansion itself is, of course, infinite, at each order we are assured that the number of terms to be retained is finite and well defined.
II.2 Chiral truncation error
Crucial to chiral EFT is the trucation error, defined below If observable $X$ is known at order $n$ and at order $n+1$, a reasonable estimate of the truncation error at order $n$ can be expressed as the difference between the value at order $n$ and the one at the next order:
$\Delta X_{n}=|X_{n+1}-X_{n}|\;,$ (5)
since this is a measure for what has been neglected at order $n$. To estimate the uncertainty at the highest order that we consider, we follow the prescription of Ref. Epelbaum et al. (2015). For an observable $X$ that depends on the typical momentum of the system under consideration, $p$, one defines $Q$ as the largest between $\frac{p}{\Lambda_{b}}$ and $\frac{m_{\pi}}{\Lambda_{b}}$, where $\Lambda_{b}$ is the breakdown scale of the chiral EFT, for which we assume 600 MeV. The uncertainty of the value of $X$ at N3LO is then given by:
$\Delta X=\max\{Q^{5}|X_{LO}|,Q^{3}|X_{LO}-X_{NLO}|,Q^{2}|X_{NLO}-X_{N^{2}LO}|,$
$Q|X_{N^{2}LO}-X_{N^{3}LO}|\}\;,$ (6)
where $p$ could be identified with the Fermi momentum at the density under consideration.
@@ II.2 Chiral truncation error
We close this section with a warning. While the predictions at N2LO are fully ab initio, a warning is in place for current N3LO calculations. As pointed out in Ref. Epelbaum et al. (2020), there is a problem with the regularized three-nucleon-force (3NF) at N3LO (and higher orders): in all present nuclear structure calculations. The N3LO 3NFs currently in use are regularized by a multiplicative regulator applied to the 3NF expressions as derived from dimensional regularization. This approach leads to violation of chiral symmetry at N3LO and destroys the consistency between two- and three-nucleon forces Epelbaum et al. (2020, 2023). Consequently, none of the current calculations that include 3NFs at N3LO (and beyond) can be considered truly ab initio. An appropriate symmetry-preserving regulator Epelbaum et al. (2020) should be applied to the 3NF at N3LO from Refs. Bernard et al. (2008, 2011). At the present time, reliable predictions exist only at N2LO, NLO, and LO. However, for the few fully ab initio calculations, the precision at N2LO is unsatisfactory. A first step towards deriving consistently regularized nuclear interactions in chiral EFT has been proposed in Refs. Krebs and Epelbaum (2023a, b). It requires the cutoff to be introduced already at the level of the effective Lagrangian. A path integral approach Krebs and Epelbaum (2023a) can then be applied to the regularized chiral Lagrangian to derive nuclear forces through the standard power counting of chiral EFT. We will come back to this point in our conclusions and outlook.
@@ III.1 Subluminar and non-conformal
The content of this section is based on Ref. Sammarruca and Ajagbonna (2025), where we discussed how causality and maximum-mass constraints pose considerable restrictions on the general features of the high-density EoS, if such continuation follows an ab intio portion.
Parametrizations of the high-density EoS in terms of the speed of sound can be used as an alternative to, or in combination with the popular polytropic extensions. Defining $\rho_{0}$, $\epsilon_{0}$, and $P_{0}$ as the baryon density, the energy density, and the pressure at threshold (the density at which the EoS parametrization is attached to the previous piece), we write Kanakis-Pegios et al. (2021); Tews et al. (2018), for $i\geq 1$,
@@
where the constants $c_{1}$ and $c_{2}$ are determined from continuity of the speed of sound and its derivative at the threshold density. We find this to be a practical choice for a function that should approach a limit asymptotically. The main point is to preserve causality, which could be done, for instance, by attaching additional polytropes with appropriate steepness so as to control maximum mass and causality.
@@
We found Sammarruca and Ajagbonna (2025) that a better ansatz, for the purpose of achieving high maximum masses while respecting causality at any density, is to first continue the EoS with a relatively steep polytrope,
$P(\rho)=\alpha\rho^{\gamma}\;,$ (12)
and then follow with a parametrization as in Eqs. (1011), which will remain causal by construction. The matching densities are 0.277 $fm^{-3}$ and 0.563 $fm^{-3}$ Sammarruca and Ajagbonna (2025). We underlined that this scenario is fundamentally related to the softness of the chiral predictions. Therefore, the nature of the predictions at normal density have a far-reaching impact, which extends to densities up to a few times normal density.
That being said, piecewise polytropic extensions are quite general and remain a popular extension method. We found that, leveraging the flexibility offered by Eq.(12) over a limited high-density interval, would allow us to control the maximum mass, whereas the causality-preserving pressure-density relation is a more practical tool (than piecewise polytropes) to reach super-high densities.
Table 1: Some neutron star properties corresponding to the red and the blue M(R) relations shown on the left-hand side of Fig. 2. The Table is taken from Ref. Sammarruca and Ajagbonna (2025).
@@ III.2 Subluminar and conformal
In the QCD limit of deconfined quarks in the presence of asymptotic freedom, quarks should behave like free fermions. Some perturbative QCD calculations Brandes et al. (2023) support the conformal limit, $(v_{s}/c)^{2}=1/3$. We observed that an EoS that’s subconformal at all densities (that is, Eq. (10) with the asymptotic limit replaced by 1/3), cannot generate sufficiently large masses. On the other hand, the speed of sound can be asymptotically conformal and non-monotone. A scenario where the speed of sound peaks around few to several times nuclear density and then falls back to approach the QCD limit for deconfined quark matter would signify some sort of phase transition with the conformal limit reached well beyond central densities of the heaviest observed neutron stars. The superconformality condition, $\Big(\frac{v_{s}}{c}\Big)^{2}>1/3$, satisfied on the average in neutron stars may have fundamental implications on the trace of the energy-momentum tensor at the center of rotating neutron stars Mendes et al. (2024).
Here, we will explore parametrizations where the speed of sound is consistent with the conformal limit as the density approaches infinity Tews et al. (2018). As an extension of Eq. (10), a natural choice might be to add a second, skewed, gaussian function, which approaches the asymptotic value of 1/3:
@@
We solve for $c_{1}$ and $c_{2}$ as done above, matching Eq. (13) and its derivative to the values at threshold, within a chosen parameter space for the remaining three coefficients, $c_{3},c_{4},c_{5}$. The widths $w$ and $w_{2}$ are obtained by normalizing both gaussians in the standard way. Although this procedure can provide a very large number of possibilities, we observed that the roots, $c_{1}$ and $c_{2}$, converge easily within limited ranges for the free parameters: $c_{3}\approx[0,1]$, $c_{4}\approx[1,4]fm^{-3}$, $c_{5}\approx[-50,50]$. In fact, the roots tend to cluster around a few pairs of $c_{1},c_{2}$ values, weakly sensitive to the large variations in $c_{5}$, and only moderately sensitive to variations of $c_{3}$ and $c_{4}$. As a consequence, those parametrizations produce very similar M(R) results. A representative case is shown by the green curves on the left-hand side of Fig. 2, where we have chosen a polytropic index of 3.5 to ensure that the new curves would stand out in between the other two sets from. The corresponding curves for the dimensionless speed of sound are seen to approach the conformal limit on the right-hand side of Fig. 2. We also note that implementing the conformal limit at all densities would not allow sufficiently large masses. Maximum masses as large as those recently observed require a rapid rise of the pressure soon after the first matching point.
It is interesting to observe that, on the left panel of Fig. 2, the green curves are what one would expect from the non-conformal calculation and an adiabatic index that’s in between the other two. In contrast, the green predictions on the right are dramatically different from the red or blue predictions. This suggests that the maximum mass is only weakly sensitive to the super-high density behavior of the EoS.
Figure 2: Left: the red and blue M(R) curves are as described in Ref. Sammarruca and Ajagbonna (2025). Green curves: M(R) relations at N3LO (dashed) and at N2LO (solid) where the polytropic part of the EoS has adiabatic index equal to 3.5, and the last continuation is based on the conformal speed of sound parametrization, Eq. (13). Right: Dimensionless speed of sound squared corresponding to each of the cases displayed on the left. See text for additional explanation.